my room has an area of 25feet and 14yards. what is the area of the classroom
Answer by jim_thompson5910(35256)

1 yard = 3 ft


14*1 yards = 14*3 ft


14 yards = 42 ft


-------------------------------------------------------


The classroom is 25 feet by 42 feet


The area is 25*42 = 1,050 square feet